OpenStudy (anonymous):

Solve by completing the square. 2x^2-12x+26=0

OpenStudy (anonymous):

I am sorry it's 2 (x^2-6x) +26 = 0 2 (x^2 - 6x + 9) + 8 = 0 2(x^2-3)^2 + 8 = 0